OpenStudy (anonymous):

suppose 8% of new computers malfunction within the first year. A school buys 15, what's the probability that more than 2 will malfunction in the first year

OpenStudy (kropot72):

You need to find the probabilities of 0, 1 and 2 malfunctions. Then these three values of probability are added, and the total subtracted from 1 to find the probability that more than 2 will malfunction in the first year. Do you want me to give help?

OpenStudy (anonymous):

yes please

OpenStudy (kropot72):

The probability that a computer will not malfunction is 1.00 - 0.08 = 0.92. The probability that 0 will malfunction is therefore: \[P(0\ malfunction)=(0.92)^{15}=you\ can\ calculate\] Can you do this calculation and post the result?

OpenStudy (anonymous):

0.92^15 = 0.28629

OpenStudy (kropot72):

Good. Rounded up it is P(0 malfunction) = 0.2863. Using the binomial distribution the probability that 1 will malfunction is: \[P(1\ malfunction)=\left(\begin{matrix}15 \\ 1\end{matrix}\right)\times (0.08)^{1} \times (0..92)^{14}=you\ can\ calculate \]

OpenStudy (kropot72):

Sorry for the typo. The calculation can be written as: \[P(1\ malfunction)=15 \times0.08\times(0.92)^{14}=you\ can\ calculate\]

OpenStudy (kropot72):

Using the binomial distribution the probability that 2 will malfunction is: \[P(2\ malfunction)=\left(\begin{matrix}15 \\ 2\end{matrix}\right) \times (0.08)^{2} \times (0.92)^{13}=105\times(0.08)^{2} \times (0.92)^{13}=you\ can\ calculate\]

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)